40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 26 17:45:21 2014, Steve, Update, VAC, vac pressure rose to 1.3 mTorr noN2.pngbacktoVacNormal.png
    Reply  Sun Dec 28 07:29:16 2014, Steve, Update, VAC, vacuum is normal 3d.png
    Reply  Tue Feb 3 15:20:13 2015, Steve, Koji, Update, VAC, Pneumatic pressure is being read again N2pressureReadingIsBack.png
    Reply  Sun Mar 22 18:42:40 2015, Steve, Update, VAC, vac pressure rose to 1.2mTorr NoN2.pngPressureReadingWorks.png
       Reply  Mon Mar 23 10:36:55 2015, ericq, Update, VAC, Pressure watch script 
          Reply  Fri Apr 24 17:30:32 2015, Jenne, Update, VAC, Pressure watch script broken 
             Reply  Sat Apr 25 18:50:47 2015, ericq, Update, VAC, Pressure watch script broken 
                Reply  Sat Apr 25 22:17:49 2015, rana, Update, CDS, MXstream restart script working (beta) 
                   Reply  Tue May 19 17:03:34 2015, Koji, Update, CDS, MXstream restart script working (beta) 
                      Reply  Tue May 19 18:38:33 2015, rana, Update, CDS, MXstream restart script working (beta) 
          Reply  Mon Apr 27 14:37:55 2015, Steve, Update, VAC, N2 pneumatic pressure watch  
             Reply  Thu Apr 30 16:30:25 2015, Steve, Update, VAC, N2 pneumatic pressure watch set up 
          Reply  Tue Jan 31 14:14:29 2017, rana, Update, VAC, Pressure watch script 
       Reply  Sun May 15 19:48:08 2016, Steve, Update, VAC, run out of N2 noN2.png
          Reply  Tue Jun 14 10:13:34 2016, Steve, Update, VAC, temporary N2 supply line afterN2work.png
             Reply  Fri Jun 17 15:22:31 2016, gautam, Update, VAC, N2 supply line restored after retiling c0vac_06172016.png
Message ID: 10843     Entry time: Fri Dec 26 17:45:21 2014     Reply to this: 10845   10968   11156
Author: Steve 
Type: Update 
Category: VAC 
Subject: vac pressure rose to 1.3 mTorr 

We run out of N2 for the vacuum system. The pressure peaked at 1.3 mTorr with MC locked. V1 did not closed because the N2 pressure sensor failed.

We are back to vac normal. I will be here tomorrow to check on things.

Attachment 1: noN2.png  287 kB  | Hide | Hide all
noN2.png
Attachment 2: backtoVacNormal.png  293 kB  | Hide | Hide all
backtoVacNormal.png
ELOG V3.1.3-